(1 point) Find y as a function of t if 250y + 4y + 3y = 0, y(0) = 3. y(0) = 6, y(t) = Note: This problem cannot interpret
Answer #1

2509 +44 + 3y z 0 - 0 For Complementary function [C. F) 250y + 4y + 3y zo (250D + 4D + 3) y zo, D=1 250D² + 4D +3=0 D = - 4At, yo = 3 - - - 9 + entzun = 3 = -6 +62 1746 3 [c2 = 27. go 46 1746 23 250 [ab] Henee Y t = ens [6 68 [1746 )t + 27.90 Sin